# Fino1: On the Transferability of Reasoning-Enhanced LLMs to Finance

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have shown strong general reasoning abilities, yet their effectiveness in financial reasoning remains underexplored. In this study, we comprehensively evaluate 16 powerful reasoning and general LLMs on three complex financial tasks involving financial text, tabular data, and equations, assessing numerical reasoning, tabular interpretation, financial terminology comprehension, long-context processing, and equation-based problem solving. Our results show that while better datasets and pretraining improve financial reasoning, general enhancements like CoT fine-tuning do not always yield consistent gains. Moreover, all reasoning strategies face challenges in improving performance on long-context and multi-table tasks. To address these limitations, we develop a financial reasoning-enhanced model based on Llama-3.1-8B-Instruct, by CoT fine-tuning and reinforcement learning with domain-specific reasoning paths. Even with simple fine-tuning with one financial dataset, our model achieves a consistent 10% performance improvement across tasks, surpassing all 8B models and even Llama3-70B-Instruct and Llama3.1-70B-Instruct on average. Our results highlight the need for domain-specific adaptations in financial tasks, emphasizing future directions such as multi-table reasoning, long-context processing, and financial terminology comprehension. All our datasets, models [^1] https://huggingface.co/TheFinAI, and codes [^2] https://github.com/The-FinAI/Fino1 are publicly available. Furthermore, we introduce a leaderboard for benchmarking future datasets and models [^3] https://huggingface.co/spaces/TheFinAI/open-finllm-reasoning-leaderboard.
@@ 1 Introduction
Financial tasks inherently demand rigorous reasoning, requiring not only numerical computations but also a deep contextual understanding of financial terminology, regulations, and economic principles Xie et al. (2023, 2024a). Unlike general reasoning tasks, financial reasoning necessitates (1) comprehension of domain-specific terminology, (2) the ability to establish and manipulate relationships between financial concepts and numbers to derive accurate conclusions, and (3) an understanding of the structure and content of financial tables Xie et al. (2024b). Given the specialized nature of financial decision-making, assessing the reasoning capabilities of LLMs in this domain is crucial for bridging the gap between general AI reasoning and financial applications. Effective financial reasoning requires models to integrate textual, numerical, and structured data, highlighting the need for domain-specific adaptations to enhance AI-driven financial analysis.
@@
In this study, we aim to comprehensively evaluate the performance of existing powerful reasoning models and conduct an in-depth analysis of their capabilities in financial tasks. Through this evaluation, we aim to provide insights into the strengths and limitations of existing reasoning models in financial applications, informing future advancements in financial AI research. We focus on the following key aspects: (1) Transferability of general-domain reasoning enhancements – Assessing whether reasoning-enhanced LLMs, originally optimized for general-domain reasoning, also demonstrate improved performance in financial reasoning tasks. (2) Impact across financial tasks – Analyzing how the reasoning enhancement process influences LLM performance across different types of financial tasks. (3) Scaling and performance gap – Investigating the differences between smaller reasoning models and large-scale LLMs in financial reasoning. (4) Limits of LLMs in complex financial tasks – Evaluating the extent to which LLMs can handle extremely complex financial reasoning challenges.
We evaluated the performances of 16 large language models from the GPT Achiam et al. (2023), LLaMA Dubey et al. (2024), DeepSeek Liu et al. (2024), and Qwen Yang et al. (2024) families on three financial datasets, FinQA Chen et al. (2021), DM-Simplong Zhao et al. (2024), and XBRL-Math Wang et al. (2025), which incorporate financial text, tabular data, and equations as inputs. Our selection includes both general purpose and reasoning-enhanced models, such as DeepSeek-V3 Liu et al. (2024) and DeepSeek-R1 Guo et al. (2025), spanning from smaller models like LLaMA 3-8B-Instruct Dubey et al. (2024) to high-capacity models like GPT-o1 Jaech et al. (2024), ensuring a comprehensive assessment across different computational scales. These three financial reasoning datasets are designed to evaluate distinct aspects of financial understanding. Collectively, they assess models’ abilities in numerical reasoning, tabular data interpretation, financial terminology comprehension, long-context processing, and equation-based problem solving. Through this evaluation, our objective is to examine the capacity of LLMs to process financial concepts, reason over structured and unstructured financial data, and perform complex mathematical operations, offering deeper insight into their strengths and limitations in financial analysis tasks.
From experimental results and analysis, we have the following findings:
(1) General reasoning enhancement strategies, such as those employed in DeepSeek-R1 and GPT-o1, do not consistently improve performance in financial domain tasks. In fact, these models exhibit a performance decline compared to general-purpose models like DeepSeek-V3 and GPT-4o, suggesting challenges in understanding financial terminology and reasoning over long contexts and multi-table data.
(2) While improvements in pretraining methods and increased training data in general domains, as well as increasing model size all contributes to performance gains.For financial tasks, model performance tends to plateau at an upper boundary. Specifically, after reaching 70B or even 32B parameters, financial reasoning tasks show no significant improvement.
(3) The performance of different reasoning enhancement strategies in financial tasks can vary significantly. For instance, general-domain techniques like Chain-of-Thought and reflection-tuning in GPT-o1, designed for math and code, show no improvements in financial reasoning. In contrast, results show that DeepSeek and Qwen-Math incorporate enhance formula computations and structured data parsing using tool-integrated reasoning (TIR) and the process reward model (PRM), and knowledge distilled from general-purpose models. While these methods improve numerical accuracy in financial reports and market data, their impact on financial terminology comprehension, long-context modeling, and multi-table inference remains limited.
Based on the evaluation results, we developed a reasoning-enhanced language model specifically designed for financial tasks. We selected Llama-3.1-8B-Instruct Dubey et al. (2024) as the backbone model due to its superior performance among small-scale LLMs. To enhance its reasoning capabilities, we employed Chain-of-Thought (CoT) fine-tuning and reinforcement learning (RL), inspired by the four backtracking strategies from HuatuoGPT-o1 Chen et al. (2024) and the FinQA training dataset. To construct effective reasoning paths for fine-tuning, we leveraged GPT-4o, the best-performing closed-source model, as it achieved comparable results across all datasets. Experimental results indicate that even a simple fine-tuning approach using the reasoning paths from a single financial dataset, FinQA, led to significant performance improvements across multiple datasets. In some cases, our model’s performance approached that of advanced closed-source reasoning models, such as GPT-o3-mini and GPT-o1. These results demonstrate that reasoning enhancement tailored to financial domain tasks and datasets can significantly boost model performance in financial applications. In contrast, reasoning enhancement in general-domain tasks does not yield effective or consistent improvements for financial tasks, highlighting the necessity of domain-specific fine-tuning for optimal results.
To the best of our knowledge, we are the first to systematically explore the reasoning capabilities of enhanced LLMs in factual reasoning tasks across various financial inputs, including financial reports, tables, equations, and structured XBRL texts. Additionally, we investigate the effectiveness of reasoning enhancement strategies using financial data in adapting general-domain models for financial tasks. Overall, our results show that reasoning enhancement strategies designed for general domains do not consistently yield stable and effective improvements in financial reasoning tasks, whereas reasoning enhancement with financial data leads to significant performance gains. Furthermore, we highlight several key directions for advancing reasoning-enhanced LLMs in the financial domain, including enhancing models’ understanding of financial knowledge and terminology, improving their ability to handle multi-table reasoning and long financial contexts, and developing well-designed and diverse reasoning enhancement strategies to improve LLM performance in financial reasoning tasks.
@@ 2.1 Datasets
To evaluate the overall capabilities of existing reasoning-oriented and general LLMs in handling financial reasoning tasks, we select three tasks with different attributes:
(1) FINQA Chen et al. (2021): FinQA is a large-scale dataset designed to facilitate research in numerical reasoning within the financial domain. It consists of expert-annotated question-answer pairs and is distinctive in its focus on complex numerical reasoning, requiring the integration of both structured data (such as tables) and unstructured data (such as textual descriptions) found within financial reports. The dataset contains a vast number of financial terms, posing a challenge for models to understand domain-specific terminology and extract relevant information accurately.
(2) DocMath (simplong) Zhao et al. (2024): The DocMath-Eval dataset is a comprehensive benchmark designed to evaluate the numerical reasoning capabilities of LLMs, with a primary focus on long contexts consisit of specialized documents and tables. In particular, the simpllong subset is specifically designed to evaluate LLMs’ numerical reasoning abilities over long financial or specialized documents that contain multiple and multi-tiered tables. It challenges models to extract relevant information from extended contexts and perform reasoning across complex tabular structures.
(3) XBRL-Math Wang et al. (2025): This dataset is designed to evaluate the numerical reasoning capabilities of LLMs in the context of XBRL (eXtensible Business Reporting Language) filings. Different from regular financial reports, XBRL files are well structured that encode financial data using standardized taxonomies, ensuring consistency and interoperability across regulatory filings. It contains structured financial documents with US GAAP XBRL tags, equations, and multi-tiered numerical relationships. This dataset challenges models to extract, interpret, and reason across interconnected financial terms and formulas, such as APR = ((Fees + Interest) / Principal) × (365 / Days in Loan Term). Understanding XBRL taxonomy, hierarchical disclosures, and numerical dependencies is essential for accurate financial data analysis, making this dataset a benchmark for evaluating LLMs’ ability to handle complex financial reasoning tasks.
@@ 2.2 Evaluated Models
To conduct a comprehensive evaluation of existing LLMs, we select models from the GPT, LLaMA, DeepSeek, and Qwen families, which have demonstrated effectiveness in various natural language processing tasks. Our selection includes models of different scales, ranging from small-sized models (8B) to large models (70B), as well as even larger ones like DeepSeek-R1 (approximately 700B). Additionally, we incorporate the most advanced closed-source models, such as GPT-4o and GPT-o1. Finally, all the models we use are listed following: (1) GPT-4o Hurst et al. (2024): OpenAI’s latest flagship model, featuring multimodal capabilities, improved efficiency, and real-time reasoning across text, image, and audio modalities. (2) GPT-o1 Jaech et al. (2024): A highly optimized variant within OpenAI’s GPT series, designed for enhanced reasoning and problem-solving with strong step-by-step thought processes. (3) GPT-o3-mini: A compact version of the o3 model, optimized for efficiency while maintaining robust language processing abilities. (4) DeepSeek-V3 Liu et al. (2024): DeepSeek’s Mixture-of-Experts (MoE) model with 671 billion parameters, activating 37 billion per token for efficient reasoning tasks. (5) DeepSeek-R1 Guo et al. (2025): DeepSeek’s first-generation reasoning model, achieving performance comparable to OpenAI’s o1 across math, code, and reasoning tasks. (6) Qwen2.5-72B-Instruct Yang et al. (2024): Alibaba’s open-source model demonstrating superior performance in Chinese and English tasks with advanced instruction tuning. (7) Qwen2.5-72B-Instruct-Math Yang et al. (2024): A specialized variant fine-tuned for mathematical reasoning, achieving strong results in complex problem-solving tasks. (8) Llama-series Models Dubey et al. (2024): This is the most widely adopted open-source large language model family, setting industry benchmarks for accessibility, efficiency, and performance. These models are trained on extensive datasets incorporating diverse multilingual and domain-specific data, leveraging high-quality pretraining and instruction tuning methodologies. In our experiments, we mainly focus on instruction finetune versions of these models ranging from 8B to 70B. (8) DeepSeek-R1-Distill Models Guo et al. (2025): This series of models is supervised fine-tuned on data distilled from DeepSeek-R1, leveraging its reasoning capabilities, with backbone models selected from Llama3.3-70B-Instruct, Qwen 2.5-32B, Qwen 2.5-14B, and Llama3.1-8B to balance efficiency and performance in instruction-following tasks.
Model Name Parameters Reasoning Enhanced Context Window Size Close/Open Source Reasoning Enhanced Training Data
@@
GPT-o1 Unknown Yes 128k Closed Public and properity data (Human-annotated CoT, MCTS-assisted Synthetic data)
GPT-o3-mini Unknown Yes 128k Closed Public and properity data
@@ 2.3 Evaluation Settings
In each dataset, the LLMs are presented with a context consisting of financial reports, financial tables, equations, and mathematical symbols, along with a question. Insired by previous work Zhao et al. (2024), a generic prompt is used to instruct the model to answer the question based on the provided context. The models are required to extract relevant information and perform mathematical computations using numbers and concepts derived from the context. This necessitates strong logical reasoning abilities within financial contexts. The results are typically presented in a mathematical format, but due to variations in numerical representation—such as percentages and different rounding strategies—the final outputs can differ. Furthermore, reasoning-focused models, such as DeepSeek-R1 and GPT-o1, tend to provide a step-by-step reasoning process before arriving at the final answer. Inspired by DocMath, we employ an LLM-as-judge approach to evaluate model performance. First, an LLM (GPT-3.5-Turbo) is used to extract the final answer from the model’s output. Then, a mathematical comparison method is applied to determine whether the result is correct.
@@
For large models like DeepSeek-V3 and DeepSeek-R1, or closed-source models like the GPT series, we conduct our experiments using APIs provided by togetherAI [^4] https://www.together.ai and OpenAI [^5] https://platform.openai.com. Specifically, we utilize GPT-o1-2024-12-17, GPT-o3-mini-2025-01-31, and GPT-4o-2024-08-06 for our evaluations. For other models with 72B parameters or fewer, we generate results using the VLLM framework on 4 A100 GPUs (80GB memory per GPU), with a maximum generation length of 512 tokens.
2.4 Fino1: LLMs for Financial Reasoning
To gain a deeper insight into how domain-specific reasoning enhancement benefits financial tasks compared to general reasoning enhancement, we developed Fino1, a reasoning-enhanced LLM specifically designed for financial tasks.
WTo enhance the domain-specific reasoning abilities of LLMs, we first construct high-quality training data with domain-specific reasoning paths based on the FinQA dataset. FinQA consists of over 5,000 open-ended question-answer pairs derived from financial reports and tables. This comprehensive dataset integrates both textual and tabular financial data, presenting a moderate level of difficulty, making it well-suited for training reasoning-based models.

### v2

@@ Abstract
While large language models (LLMs) have shown strong general reasoning capabilities, their effectiveness in financial reasoning—which is crucial for real-world financial applications—remains underexplored. In this study, we conduct a comprehensive evaluation of 24 state-of-the-art general and reasoning-focused LLMs across four complex financial reasoning tasks involving financial text, tabular data, and equations. We assess key capabilities such as numerical reasoning, tabular interpretation, financial terminology comprehension, long-context understanding, and equation-based problem solving. Our analysis reveals that while data quality and pretraining contribute to performance, general techniques like chain-of-thought (CoT) fine-tuning offer limited gains in financial tasks. To address this, we propose two domain-adapted models, Fino1-8B and Fino1-14B, trained with CoT fine-tuning and reinforcement learning using domain-specific reasoning paths. Our models are trained on a carefully curated dataset integrating high-quality examples from diverse sources, covering financial reports, tables, equations, and structured XBRL texts. Despite limited training data, they achieve an 7–9% performance improvement, outperforming several advanced LLMs—including GPT-o1, GPT-o3-mini, GPT-4.5, and comparable with DeepSeek models (V3 and R1)—demonstrating strong practical value in resource-constrained scenarios. Our findings highlight the need for domain-specific adaptations in financial reasoning, and we release all datasets, models, and code for future research [^1] https://github.com/The-FinAI/Fino1.
@@ 1 Introduction
Financial reasoning is essential in the financial domain, where professionals must extract key insights from lengthy reports and multi-table documents. It requires not only numerical computation but also a deep understanding of financial terminology, regulations, and economic concepts Xie et al. (2023, 2024a). Unlike general reasoning tasks, financial reasoning involves (1) domain-specific language comprehension, (2) reasoning over financial concepts and numbers, and (3) interpreting structured data such as financial tables Xie et al. (2024b). Evaluating LLMs in this context is critical for closing the gap between general AI reasoning and real-world financial applications.
@@
To address this gap, we firstly comprehensively evaluate the performance of existing powerful reasoning models and conduct an in-depth analysis of their capabilities in financial tasks, to provide insights into the strengths and limitations of existing reasoning models in financial applications. We focus on the following key aspects: (1) Transferability of general-domain reasoning enhancements – Assessing whether reasoning-enhanced LLMs, originally optimized for general-domain reasoning, also demonstrate improved performance in financial reasoning tasks. (2) Impact across financial tasks – Analyzing how the reasoning enhancement process influences LLM performance across different types of financial tasks. (3) Scaling and performance gap – Investigating the differences between smaller reasoning models and large-scale LLMs in financial reasoning. (4) Limits of LLMs in complex financial tasks – Evaluating the extent to which LLMs can handle extremely complex financial reasoning challenges.
We evaluated 24 LLMs from the GPT Achiam et al. (2023), LLaMA Dubey et al. (2024), DeepSeek Liu et al. (2024), and Qwen Yang et al. (2024) families on four financial reasoning benchmarks: FinQA Chen et al. (2021), DM-Simplong Zhao et al. (2024a), XBRL-Math Wang et al. (2025) and DM-Complong. These datasets span financial texts, tables, and equations, enabling evaluation of models in numerical reasoning, domain-specific terminology comprehension, long-context processing, and equation-based reasoning tasks. Our evaluation encompasses both general-purpose and reasoning-enhanced models, ranging from smaller-scale models (e.g., LLaMA 3-8B-InstructDubey et al. (2024)) to high-capacity models (e.g., GPT-o1Jaech et al. (2024)), as well as reasoning-augmented models incorporating strategies such as data sampling and financial Chain-of-Thought (CoT) construction. Experimental results reveal three key insights: (1) Reasoning-enhanced models like DeepSeek-R1 and GPT-o1 do not consistently outperform general-purpose models on financial tasks, likely due to challenges related to domain-specific terminology and multi-table reasoning; (2) Performance gains from larger model sizes and improved pretraining tend to plateau beyond 32B or 70B parameters;
To develop reasoning models tailored for financial domain, we build novel financial reasoning-enhanced large language models Fino1-8B and Fino1-14B based on LLaMA-3.1-8B-InstructDubey et al. (2024) and Qwen2.5-14B-InstructYang et al. (2024), chosen for their strong performance among smaller-scale LLMs. To enhance reasoning capabilities, we apply CoT fine-tuning and reinforcement learning (RL), inspired by the training strategies of HuatuoGPT-o1 Chen et al. (2024). We leverage a diverse set of high quality financial reasoning datasets—including FinQA Chen et al. (2022), TATQA Zhu et al. (2021), DocMath (Simpshort and Compshort) Zhao et al. (2024b) , DocFinQA Reddy et al. (2024a), and BizBench-QA Krumdick et al. (2024)—covering tasks such as long-context question answering, complex reasoning over financial documents, and table-text integration. We construct high-quality reasoning paths using GPT-4o, combined with a novel data filtering process that selects more challenging samples for training—particularly for the 14B model—ensuring the model learns from difficult, domain-specific reasoning cases rather than generic QA pairs. Experimental results show that this targeted training strategy enables our models to surpass the performance of advanced closed-source models such as GPT-o3-mini and GPT-o1, and approach most powerful LLMs like DeepSeek-V3, DeepSeek-R1, and GPT-4o, demonstrating that domain-specific reasoning-enhanced training with limited, high-quality data can be highly effective.
To the best of our knowledge, this work presents the first comprehensive study of reasoning-enhanced LLMs for financial tasks involving diverse inputs such as financial documents, tables, equations, and structured XBRL texts. First, we conduct a systematic evaluation of 24 general-purpose and reasoning-enhanced LLMs across multiple financial reasoning benchmarks. Second, we introduce Fino1, the first suite of domain-specific, reasoning-enhanced LLMs, trained using GPT-4o-distilled reasoning paths with a novel data curation and filtering strategy. Third, our results show that domain-specific reasoning training with limited high-quality data consistently outperforms general-domain strategies and even surpasses larger closed-source models. We release all models, datasets, and code to support future research in financial AI.
@@ 2.1 Datasets
To evaluate the financial reasoning capabilities of reasoning-oriented and general LLMs, we select four diverse benchmark tasks. (1) FINQA Chen et al. (2021) is a large-scale dataset that emphasizes complex numerical reasoning in financial reports, requiring the integration of structured tables and unstructured text, along with a deep understanding of domain-specific terminology. (2) DocMath (Simplong, Complong) Zhao et al. (2024a) evaluates LLMs’ numerical reasoning abilities over long, specialized documents with multi-tiered tables, testing their capacity to extract and reason over extended contexts. The simplong subset contains relatively easier questions, while the complong subset includes more complex questions that require advanced reasoning across tables and context. (3) XBRL-Math Wang et al. (2025) targets reasoning over structured financial data in XBRL filings, which involve standardized taxonomies, hierarchical disclosures, and formulas, challenging models to interpret intricate numerical relationships.
@@ 2.2 Evaluated Models
[…]
The evaluated models are summarized as follows: GPT-4o Hurst et al. (2024): OpenAI’s flagship model with multimodal capabilities and improved real-time reasoning. GPT-o1 Jaech et al. (2024): A reasoning-optimized variant in the GPT series with strong step-by-step problem-solving ability. GPT-o3-mini: A compact version of GPT-o3, optimized for efficiency with robust language understanding. DeepSeek-V3 Liu et al. (2024): A 671B Mixture-of-Experts (MoE) model with 37B active parameters per token, designed for efficient reasoning. DeepSeek-R1 Guo et al. (2025): A first-generation reasoning model with competitive performance on math, code, and reasoning tasks. Qwen2.5-72B-Instruct Yang et al. (2024): A strong multilingual model with advanced instruction tuning. Qwen2.5-72B-Instruct-Math: A math-specialized variant fine-tuned for complex problem solving. LLaMA-series Dubey et al. (2024): Widely adopted open-source models (8B–70B) with strong instruction-tuned variants used in our experiments. DeepSeek-R1-Distill Guo et al. (2025): Distilled models from DeepSeek-R1, fine-tuned using LLaMA and Qwen backbones to balance performance and efficiency. Other reasoning-enhanced models: Includes S1-32BMuennighoff et al. (2025), Limo Ye et al. (2025), and FinR1-7B Liu et al. (2025b), incorporating reasoning strategies such as financial CoT construction and task-specific fine-tuning.
Model Name Parameters Enhanced for Reasoning Tasks Context Window Size Close/Open Source Reasoning Enhanced Training Data
@@
GPT-o1 Unknown Yes 128k Closed Public and proprietary data (Human-annotated CoT, MCTS-assisted Synthetic data)
GPT-o3-mini Unknown Yes 128k Closed Public and proprietary data
@@
Limo 32B Yes 128k Open Public and proprietary data with Qwen filteration
S1-32B 32B Yes 128k Open Public and proprietary data + original data, filtered by quality, difficulty, diversity
Qwen/QwQ-32B 32B Yes 128k Open -
Qwen2.5-14B-Instruct 14B No 128k Open -
@@
Qwen2.5-7B-Instruct 7B No 128k Open -
FinR1-7B 7B Yes 128k Open 60k CoT data for financial tasks
@@ 2.3 Evaluation Settings
In each dataset, LLMs are presented with a context containing financial reports, tables, equations, and symbols, along with a question. Following prior work Zhao et al. (2024a), we use a generic prompt (see appendix D for more details) to instruct models to answer based on the given context. Models must extract relevant information and perform mathematical reasoning using numbers and concepts from the context. While outputs are typically in mathematical form, variations in formats (e.g., percentages, rounding) may cause differences. Reasoning-focused models like DeepSeek-R1 and GPT-o1 often generate step-by-step reasoning. Inspired by DocMath Zhao et al. (2024a), we adopt an LLM-as-judge evaluation approach: GPT-3.5-Turbo extracts final answers, followed by a mathematical comparison to assess correctness.
@@
For large models (e.g., DeepSeek-V3, DeepSeek-R1, GPT series), we run evaluations via APIs from TogetherAI [^2] https://www.together.ai and OpenAI [^3] https://platform.openai.com, using versions GPT-o1-2024-12-17, GPT-o3-mini-2025-01-31, and GPT-4o-2024-08-06. For models up to 72B parameters, we use the VLLM framework on 8 A100 GPUs (80GB each), with a max generation length of 512 tokens.
2.3.1 Reasoning Data Curation
To enhance domain-specific reasoning capabilities in LLMs, we curate a high-quality training dataset by integrating diverse financial reasoning benchmarks. Our dataset combines a wide range of reasoning tasks—covering numerical computation, table-text integration, cross-table correlation, and long-context understanding—reflecting the complexity of real-world financial documents.
Specifically, we incorporate data from several well-established benchmarks, including FinQA Chen et al. (2022), DocFinQA Reddy et al. (2024b), TAT-QA Zhu et al. (2021), DocMath-Eval Zhao et al. (2024b), and BizBench-QA Krumdick et al. (2024). These datasets collectively provide rich and diverse coverage across task types and difficulty levels, supporting the development of robust financial reasoning models. FinQA provides QA pairs combining textual and tabular financial data; DocFinQA extends this to long-context documents; TAT-QA focuses on complex numerical reasoning; DocMath-Eval (CompLong and SimpShort subsets) targets reasoning over multi-table and single-table scenarios; and BizBench-QA enriches the corpus with code-based financial reasoning tasks. Please see Appendix B for detailed descriptions of each dataset.
2.4 Data Collection and Reasoning Path Construction
To ensure a balanced dataset of both short and long context samples, we collect all data from the FinQA train set, TATQA, DocMath-CompLong, and DocMath-SimpShort. Additionally, we sample 1,000 examples each from BizBench-QA and DocFinQA to balance short and long reasoning tasks. In total, we gather 9,851 financial QA samples for reasoning path construction. Inspired by Ye et al. (2025); Muennighoff et al. (2025), which suggest that simple QA pairs may not effectively improve reasoning capabilities, we apply a filtering strategy to retain more challenging examples. Specifically, we use Qwen-14B-Instruct to generate answers for each question and employ GPT-3.5-Turbo as a verifier. QA pairs correctly answered by Qwen-14B are removed, resulting in a more difficult training subset of 3,472 QA pairs for optimizing more advanced models.
2.5 Reasoning Path Generation and Refinement